The force acting on a beam was measured under the same operating conditions and a sample of 8 data points was collected. This process was repeated by 9 observers. Determine the number of degrees of freedom for the pooled 9 data sets.
Answer #1

Please rate the answer. Thank you.

Solution:

The pooled degrees of freedom are given by,

DF=N-k

Where, N = Total no. of data sets collected by all observers cumulatively

k = Total no. of observers

Thus, DF = 8 × 9 - 9 = 63

Thus total no of degrees of freedom for the pooled data sets are 63.